OpenStudy (samigupta8):

For b>0 let A1 be the area bounded by x=0,x+y=1,y=bx^2 and A2 be the area bounded by y=0,x+y=1,y=bx^2 such that A1:A2=11:16 then the value of b is...

OpenStudy (samigupta8):

I came up with A1 and A2 as 11/54 and 16/54 respectively... Bt don't know how to do it next .... The thing that's annoying me is how to put the limits of integration for the area bound by curve and line...

OpenStudy (samigupta8):

@michele_laino

OpenStudy (michele_laino):

we have to determine the value of the parameter \(b\)

OpenStudy (samigupta8):

I know the question... Bt not the way to do that

OpenStudy (michele_laino):

I think that we can try to compute the areas A1 and A2

OpenStudy (samigupta8):

Well i found the areas...bt not in terms of that term b

OpenStudy (michele_laino):



OpenStudy (samigupta8):

Bcoz i don't know the intersection point of the line and the curve mentioned

OpenStudy (samigupta8):

Exactly sir! I made the diagram like that only for A1

OpenStudy (michele_laino):

we can compute the x-coordinate of the intersection point, solving this quadratic equation: \[\Large \begin{gathered} x + y = 1 \Rightarrow x + b{x^2} = 1 \Rightarrow \hfill \\ \hfill \\ \Rightarrow b{x^2} + x - 1 = 0 \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

U mean that in terms of square root and all.. It would be so messy then!

OpenStudy (michele_laino):